# Double bracket vector fields on Poisson manifolds

### v1

@@ 4.3. Degeneracy structure of the metriplectic tensor $\mathcal{M}$: red zones
where the function $W$ is given in (4.16) and $U$ and $V$ can be chosen arbitrarily. Our main theorem, Theorem 3.4, is applicable whenever we exclude $\mathcal{M}$-singular points, while singular symplectic leaves—all pointlike in our case, see (4.5)—are admissible. Thus, in view of Prop. 4.2, let us define the set
@@
Let us illustrate this for the three examples highlighted in Section 4.1:
@@
For the case $(\mathfrak{sl}_{2}^{*},\kappa,\Pi_{\mathfrak{sl}_{2}^{*}})$ of Example 4.1, resulting from $U(z)=z$ and $V=0$ in (4.18), we see that $f(x,y,z)=2C(x,y,z)$, cf. (4.6). Thus, the red zone $R$, depicted in Fig. 8A, agrees precisely with the two regular symplectic leaves obtained from $C_{lin}=0$ when the origin is excluded, see Fig. 1B. Both these leaves are thus “bad” symplectic leaves.
@@
(A) The Lie Poisson manifold $\mathfrak{sl}_{2}^{*}$, Example 4.1,
@@
4.4. Geometric interpretation, green zones and red lines
@@
4.5. Double bracket vector fields as gradients in green zones
@@
Now, $\flat_{g}$ applied to any vector field $a\partial_{x}+b\partial_{y}+c\partial_{z}$ yields the 1-form $b\mathrm{d}x+a\mathrm{d}y+c\mathrm{d}z$, and since $\sharp_{\pi}\mathrm{d}x=X_{x}$ etc, we easily find the image of (4.5) under $\sharp_{\pi}\circ\flat_{g}$ to be
@@
Plugging (4.5) into (4.5), we obtain—cf. (2.3)—the three fundamental double bracket vector fields $\partial_{\mathcal{M}}x$, $\partial_{\mathcal{M}}y$, and $\partial_{\mathcal{M}}z$, respectively.
@@
Acknowledgements. Z. Ravanpak acknowledges a scholarship “Cercetare postdoctorală avansată” funded by the West University of Timişoara, Romania. The authors are grateful to Thomas Strobl for valuable discussions and helpful suggestions to improve significantly the presentation of the paper.

### v2

@@ 6. Degeneracy structure of the metriplectic tensor: red zones
where the function $W$ is given in (6.3) and $U$ and $V$ can be chosen arbitrarily. Our main theorem, Theorem 3.4, is applicable whenever we exclude $\mathcal{M}$-singular points, while singular symplectic leaves—all pointlike in our case, see (4.5)—are admissible. Thus, in view of Prop. 6.1, let us define the set
@@
Let us illustrate this for the three examples highlighted in Section LABEL:sec:4.1:
@@
For the case $(\mathfrak{sl}(2,\mathbb{R})^{*},\kappa,\Pi_{\mathfrak{sl}(2,\mathbb{R})^{*}})$ of Example 4.1, resulting from $U(z)=z$ and $V=0$ in (6.5), we see that $f(x,y,z)=2C(x,y,z)$, cf. (4.6). Thus, the red zone $R$, depicted in Fig. 8A, agrees precisely with the two regular symplectic leaves obtained from $C_{lin}=0$ when the origin is excluded, see Fig. 1B. Both these leaves are thus “bad” symplectic leaves.
@@
(A) The Lie Poisson manifold $\mathfrak{sl}(2,\mathbb{R})^{*}$, Example 4.1,
@@
7. Geometric interpretation, green zones and red lines
@@
8. Generalized double bracket vector field as gradient vector field in green zones
@@
Now, $\flat_{g}$ applied to any vector field $a\partial_{x}+b\partial_{y}+c\partial_{z}$ yields the 1-form $b\mathrm{d}x+a\mathrm{d}y+c\mathrm{d}z$, and since $\sharp_{\pi}\mathrm{d}x=X_{x}$ etc, we easily find the image of (8) under $\sharp_{\pi}\circ\flat_{g}$ to be
@@
Plugging (8) into (8), we obtain—cf. (2.6)—the three fundamental double bracket vector fields $\partial_{\mathcal{M}}x$, $\partial_{\mathcal{M}}y$, and $\partial_{\mathcal{M}}z$, respectively.
@@
Conclusions. In this paper we have generalized the double vector field construction defined originally only on semi-simple Lie algebras to general pseudo-Riemannian Poisson manifolds. The generalized double bracket vector field has some useful properties. It is tangent to all the symplectic leaves. Moreover, when restricted to such a leaf, it proves to be of gradient type. This happens with respect to a metric on the leaf which generalizes the normal metric on adjoint orbits in compact semi-simple Lie algebras.
In the case of a Poisson manifold with metric of indefinite signature, the generalization works only on regions where the metric induced on the symplectic leaves is non-degenerate. These regions are called green zones and leaves for which the induced metric is non-degenerate everywhere are called good leaves. We have provided a characterization of such regions which avoids determining the induced metric explicitly, which can become quite complicated. All this have been illustrated for a wide class of Poisson structures on $\mathbb{R}^{3}$, which we have discussed in full detail.
Acknowledgements. Z. Ravanpak acknowledges a scholarship “Cercetare postdoctorală avansată” funded by the West University of Timişoara, Romania, and the financial support from the Spanish Ministry of Science and Innovation under grants PID2022-137909NB-C22. The authors are grateful to Thomas Strobl for valuable discussions and helpful suggestions to improve significantly the presentation of the paper.